# PICT – A Differentiable, GPU-Accelerated Multi-Block PISO Solver for Simulation-Coupled Learning Tasks in Fluid Dynamics

### v1

@@ 5.2 Backward Facing Step
Building on the insights gained from the vortex street, we target a backward-facing step (BFS) scenario, a classic separated flow produced by an abrupt change in geometry [71, 72, 73, 74]. Unlike the vortex street, where periodic shedding dominates, the BFS introduces new challenges by emphasizing spatial development over long temporal evolutions. The flow evolves spatially along the downstream region, necessitating long-term accuracy and stability to reproduce the turbulent statistics consistently. The focus shifts from periodic shedding to maintaining correct statistical properties over an extended domain. This case tests the solver’s ability to handle flows characterized by separation and reattachment dynamics. Notably, while the target is to achieve accurate statistical properties, the training process remains centered on predicting the instantaneous velocity field rather than directly targeting statistical metrics [75]. This approach emphasizes the solver’s ability to sustain physical accuracy while inherently preserving statistical consistency.
The geometry of the domain contains a gap between the step and the top wall of $h=1\,\mathrm{m}$, and a total channel height of $H=5h$, as described in B.5. In the following, $U_{\text{b}}$ denotes the bulk velocity, and $\nu$ the kinematic viscosity. The expansion ratio ($ER=H/h$) and the Reynolds number ($Re=2hU_{b}/\nu$) are considered as two of the most effective factors that influence metrics like reattachment length for BFS [71, 72]. The train and test sets are created by varying the ER and Re as detailed in table 4. In line with the vortex street case, we target a corrector learning setup, where the model should correct a low resolution simulation with 4$\times$ spatial downsampling and 4$\times$ temporal downsampling to match the high-resolution reference. We use the same NN architecture as before, and a curriculum-based training approach starting with 10 unrolled steps. Training continues with 30 steps, and concludes with unrolling 40 steps. The resulting models are denoted as $\text{NN}_{30}$ and $\text{NN}_{40}$.
Dataset No. $s/h$ ER Re Sample range ($\mathrm{s}$) Unrolled steps
@@
Figure 8 presents the streamlines illustrating the flow patterns, averaged over $t=6000\Delta t$ (equivalent to $tU_{b}/h=120$) for all cases, including the reference solution, the No-Model baseline, $\text{NN}_{30}$ and $\text{NN}_{40}$. The contour maps show velocity magnitudes $\|\mathrm{\mathbf{u}}\|$. From the reference we can obtain the location of the separation point ($X_{s}=12.46$) and reattachment point ($X_{r}=15.03$). Compared to the reference, the baseline shows significant deviations, with the separation point shifted upstream to $X_{s}=10.34$ and the reattachment length of the bottom recirculation bubble drastically shortened to $X_{r}=9.27$. Schafer [76] attributed such errors to grid-induced oscillations, where the coarse grid amplifies unstable shear layers, accelerating the transition from laminar to turbulent flow. This amplified turbulence increases vertical transport, thereby reducing the primary reattachment length. Our learned models, $\text{NN}_{30}$ and $\text{NN}_{40}$, significantly improve flow predictions, closely aligning with the reference solution by accurately capturing both separation point (at $X_{S}=12.46$) and the reattachment point ($X_{R}=15.03$) with an error of less than $6.3\times 10^{-2}$.
@@
Since the statistical result is the primary objective for current task, figure 9 shows the MSE of the temporally averaged velocity compared to the reference across a sample range of initial conditions at three simulation lengths: 100, 4000, and 6000 $\Delta t$. Clearly, our methods consistently outperform the baseline approach, with the baseline model exhibiting significantly higher MSE values. At step $=6000$, the MSE of our methods ($1.744\times 10^{-4}$) is about 110 times lower that than of the baseline ($1.928\times 10^{-2}$). Despite the elevated MSE values, the baseline shows considerable fluctuations, attributed to varying levels of difficulty for different initial states. In contrast, our models demonstrate consistently stable performance, underscoring their robustness and reliability. Building on previous studies [77, 73], we also include the comparison of the wall skin-friction coefficient on the top and bottom wall, as well as the streamwise velocity profiles at different locations downstream the change of geometry.
@@ 5.3 Turbulent Channel Flow
Figure 13: A comparison of the per-frame statistics losses from eq. (15) over a long-term rollout consisting of 7500 simulation steps. In training, the CNN models sees at most 108 steps, indicated by the dashed vertical line. All tests used the same initial condition that was sampled from an uncorrelated simulation with a turbulence statistics loss close to 1.
@@
For our final learning setup we target a three dimensional scenario with a turbulent channel flow (TCF) setup. This setup contains a periodic channel with no-slip boundaries at ±$\pm y$ which is driven by a dynamic forcing $\nu\left.\partial\overline{u}/\partial y\right|_{\text{wall}}$ to prevent a loss of energy. The TCF is a well studied scenario that requires high spatial and temporal resolution as well as long simulation times for the turbulence statistics to converge [59]. In this scenario, coarse spatio-temporal resolutions quickly yield incorrect statistics, and hence it is crucial to introduce a form of numerical modeling of the unresolved scales to obtain an accurate solution without excessive runtimes. In line with previous experiments, we train an SGS model in form of a corrector $G_{\theta}$ that is tasked to estimate a correcting force $S_{\theta}$ at a low spatial resolution of $64\times 32\times 32$. For the TCF, $G_{\theta}$ receives the instantaneous velocity and the normalized wall distance $1-|y/\delta|$ as inputs, i.e., $S^{n}_{\theta}=G_{\theta}(\mathrm{\mathbf{u}}^{n},1-|y/\delta|)$. The term $1-|y/\delta|$ is added to inform the network of the grid refinement in regions near the wall. As neural network architecture for $G_{\theta}$ we employ a simple CNN with layers using 8, 64, 64, 32, 16, 8, 4, and 3 filters, each having a kernel size of $3^{3}$. Only the last layer uses a kernel size of $1^{3}$. This gives 198931 trainable parameters in total. ReLU activations are used for all but the last layer. Different from the previous scenarios, the dynamics in the TCF are highly chaotic and matching individual realizations of the flow from simulations at different resolutions does not provide a physically meaningful learning target. Hence, we instead aim for matching the turbulence statistics via the statistic loss from eq. (13), complemented with a regularization term on the generated forcing.
@@
$L=L_{stats}+\lambda_{S}\frac{1}{N}\sum_{t=0}^{N-1}||S_{\theta}^{n}||_{2}^{2}.$ (15)
We additionally constrain the forcing to the $[-2,2]$ range to stabilize early training. Aside from the loss terms guiding the model towards producing the desired turbulence statistics, it proved essential to prevent un-physical network outputs, which in our case means violating the incompressibility assumption. To ensure divergence free flow motions, we include the gradient modification from eq. (11) for $S_{\theta}$.
@@
Based on our results from section 4.3, we consider two different approaches for the training of the CNN model: in the first one, we exclude the gradients of the linear solves from the optimization (i.e. only using $J^{\text{none}}$ from eq. (8)). In the second variant, we start in the same way, but include the terms at a later stage of training for fine tuning. The network is trained in four different phases where the warm-up steps are sampled uniformly random from $[0,0],~{}[0,12],~{}[0,14],$ and $[0,96]$, respectively. The first phase consists of 6k optimization steps, while the other three include 20k steps. The warm-up is always followed by 12 steps of unrolling, for which we backpropagate through the PICT solver. We denote the result of this initial training as CNN SGS base. We then fine-tune the network with gradients from the advection linear solve, called CNN SGS Adv, for another 20k optimization steps, again with $[0,96]$ steps of warm-up. To ensure a fair comparison, we also continue training with the initial model as CNN SGS for another 20k optimization steps, but without the additional gradient terms. As can be seen from figure 13, the extra gradients from the advection solve have no clear beneficial effect in this scenario while roughly doubling the training time. This is consistent with our results from section 4.3 where a rollout of 10 steps did not benefit from the additional gradient terms. The learning task for the TCF models potentially also benefits from the fact that the network still receives feedback for every simulation step via the turbulence statistics loss, making a backpropagated transport of the gradients through the advection unnecessary.
@@
For the no SGS and SMAG variants, the mean streamwise velocity $U^{+}$ deviates noticeably from the reference, although SMAG is closer to the target $\text{Re}_{\tau}$ despite the model having almost no influence in the near wall region. When comparing second order statistics, we can observe how the non-modelled simulation and the simple turbulence model are insufficient to correctly simulate the channel flow at the chosen resolution. Our learned corrector, on the other hand, matches the target statistics very well, and the resulting $\text{Re}_{\tau}=548$ is very close to the target value of 550. Since matching the statistics used in a training loss is expected for a well trained network, we also compare the resulting turbulent energy budgets to corresponding reference values in figure 12. Despite not being trained on these quantities, our model accurately matches the reference budget terms, although slight deviations are noticeable in this evaluation, in particular close to the wall and in the dissipation term. However, it is worth noting that both baselines fail to match any of the budget terms.
As is evident from figure 13, where the per-frame errors in terms of statistics over a long rollout are shown, our model is not only about two orders of magnitude more accurate than no model and SMAG, but also stays stable for a rollout more than 50$\times$ longer than the training horizon. All versions quickly correct the statistics from the initial state, and, after a slight deterioration beyond the training horizon, successfully keep the error at a low level. In comparison, the Smagorinsky model maintains a level of accuracy that is comparable to the simulation without turbulence modeling. When observing the instantaneous velocity fields, pictured in figure 14, we can observe how the flow in the simulation with the learned corrector maintains the topological features of flows at this Reynolds number, in particular the streamwise streaks spacing is closer to the reference than the no SGS and SMAG simulations. The latter shows a overly-smooth flow close to the wall when compared to the reference.
@@ 5.4 Runtime Performance
For the TCF cases, training a corrector NN takes about 42 hours for the initial four phases combined, using a single Nvidia GTX 1080 Ti GPU. The fine-tuning with gradients from the advection linear solve takes another 26 hours, 46% of which are needed for the backwards advection solve, while fine-tuning without them takes 13 hours. For inference, we compare the runtime performance of our GPU-based solver to OpenFOAM’s PISO implementation, running on 32 cores of a Xeon Gold 5220R. These CPUs provide 34.4 TFLOPS of compute resources, in comparison to 11.34 TFLOPS available to the PICT solve. In each case, we measure the wall clock time it takes to simulate 20ETT. OpenFOAM needs 2.5h for the finely resolved simulation at $192\times 96\times 96$. Despite the higher resolution, it yields a 36% higher aggregated error in comparison to PICT with the learned SGS model, where the error is computed across all turbulence statistics. Our solver with learned SGS model takes 223s, 38.7% of which are required for the neural network. This means that the corrected simulation with PICT is ca. 40$\times$ faster than OpenFOAM, while still matching the reference statistics with a substantially higher accuracy, and running on a single consumer-grade GPU. Naturally, the learned solver is less general than OpenFOAM, but the substantial speedup and its accuracy nonetheless point to the very significant potential of solver-NN hybrids for turbulent, three-dimensional flows.
@@ 6 Conclusion
Having established the simulator’s applicability to learning tasks, we showed its efficacy in a number of learning applications. In two challenging 2D settings, we trained stable corrector networks to yield accurate solutions with low-resolution simulations over long rollouts. The learning setup was adapted to three-dimensional turbulent channel flows, enabling the corrector to recover the statistical properties of turbulence over rollout horizons exceeding the training window by orders of magnitude, without requiring direct supervision. It additionally outperforms traditional solvers like OpenFOAM both in terms of accuracy and computational performance. Together, these tests show the efficacy of using learned corrector networks in conjunction with coarse simulations to retain high fidelity behavior in the flows. The availability of the PICT solver as open source [^2] https://github.com/tum-pbs/PICT provides a powerful foundation for advancing learned turbulence modeling for the research community.
While our solver is implemented with GPU support for increased performance, it is currently limited to a single GPU. As future work, our simulator could be extended to work in multi-GPU setups, and to include multi-grid solvers for the advection and pressure systems, from which we expect further performance benefits. Overall, the presented results are indicative of our solver’s ability to address optimization tasks like initial and boundary value reconstruction, learning for control problems, and, with an extension for differentiable transformations, potentially also for tasks like shape optimization [81]. Moreover, the solver’s full differentiability enables promising opportunities in uncertainty quantification for fluid dynamics through probabilistic and diffusion-based learning frameworks [82, 83].
@@ A.2 PISO Algorithm
which is solved for the pressure $p$. This pressure is then used to compute the corrected, divergence-free velocity $\mathrm{\mathbf{u}}^{**}$ with
@@
The pressure correction, equations 18, 19 and 20, are repeated twice [5], with an additional indicating the second round of updates. The velocity of the next time-step is then $\mathrm{\mathbf{u}}^{n+1}:=\mathrm{\mathbf{u}}^{***}$.
@@ A.3 Discretization
Set of valid neighbour cells of $P$ $\in\textbf{F}$ F
Set of faces between $P$ and F $\in\textbf{f}$ f
Set of (virtual) boundary neighbour cells of $P$ $\in\textbf{B}$ B
Set of boundary-faces between $P$ and B $\in\textbf{b}$ b
@@
Using this notation to reference values in one of the matrices is less straight-forward, since the matrix entries relate to two cells. Thus, with current cell $P$ with index $i$ and neighbor $\in\textbf{F}$ F with index $j$, the entries of a matrix $C$ are referenced as:
@@
Neighbour entry of neighbour cell (its own center/diag) ${\left[C_{F}\right]_{F}}=C_{jj}$

### v2

@@ 5.2 Backward Facing Step
Building on the insights gained from the vortex street, we target a backward-facing step (BFS) scenario, a classic separated flow produced by an abrupt change in geometry [76, 77, 78, 79]. Unlike the vortex street, where periodic shedding dominates, the BFS introduces new challenges by emphasizing spatial development over long temporal evolutions. The flow evolves spatially along the downstream region, necessitating long-term accuracy and stability to reproduce the statistics of the flow consistently. The focus shifts from periodic shedding to maintaining correct statistical properties over an extended domain. This case tests the solver’s ability to handle flows characterized by separation and reattachment dynamics. Notably, while the target is to achieve accurate statistical properties, the training process remains identical to the setup of the vortex street case, which is centered on predicting the instantaneous velocity field, rather than directly targeting statistical metrics. This approach emphasizes the solver’s ability to sustain physical accuracy while inherently preserving statistical consistency.
[…]
@@
Figure 8 presents the streamlines illustrating the flow patterns, averaged over $t=6000\Delta t$ (equivalent to $tU_{b}/h=120$) for all cases, including the reference solution, the No-Model baseline, $\text{NN}_{30}$ and $\text{NN}_{40}$. The contour maps show velocity magnitudes $\|\mathrm{\mathbf{u}}\|$. From the reference we can obtain the location of the separation point ($X_{s}=12.46$) and reattachment point ($X_{r}=15.03$). Compared to the reference, the baseline shows significant deviations, with the separation point shifted upstream to $X_{s}=10.34$ and the reattachment length of the bottom recirculation bubble drastically shortened to $X_{r}=9.27$. Similar discrepancies have been reported by Schafer [80] in 3D simulations. They attributed such errors to grid-induced oscillations, where the coarse grid amplifies unstable shear layers, accelerating the transition from laminar to chaotic flow. The amplified motions increase vertical transport, thereby reducing the primary reattachment length. Although the current case is in 2D and thus lacks the spanwise dynamics inherent to 3D flows, a comparable effect is observed: insufficient resolution leads to a significantly shortened reattachment length. This systematic deviation highlights the limitations of the baseline and the opportunity for learned models to recover more physically accurate behavior. Our learned models, $\text{NN}_{30}$ and $\text{NN}_{40}$, significantly improve flow predictions, closely aligning with the reference solution by accurately capturing both the separation point (at $X_{S}=12.46$) and the reattachment point ($X_{R}=15.03$) with an error of less than $6.3\times 10^{-2}$.
@@
Since the statistical result is the primary objective for current task, figure 9 shows the MSE of the temporally averaged velocity compared to the reference across a sample range of initial conditions at three simulation lengths: 100, 4000, and 6000 $\Delta t$. Clearly, our methods consistently outperform the baseline approach, with the baseline model exhibiting significantly higher MSE values. At step =6000=6000, the MSE of our methods ($1.744\times 10^{-4}$) is about 110 times lower that than of the baseline ($1.928\times 10^{-2}$). Despite the elevated MSE values, the baseline shows considerable fluctuations, attributed to varying levels of difficulty for different initial states. In contrast, our models demonstrate consistently stable performance, underscoring their robustness and reliability. Building on previous studies [81, 78], we also include the comparison of the wall skin-friction coefficient on the top and bottom wall, as well as the streamwise velocity profiles at different locations downstream the change of geometry.
@@ 5.3 Turbulent Channel Flow
Figure 13: A comparison of the per-frame statistics losses from eq. (16) over a long-term rollout consisting of 20000 simulation steps. In training, the CNN models sees at most 108 steps, indicated by the dashed vertical line. All cases are evaluated for 6 different initial conditions that are sampled from an uncorrelated simulation with a turbulence statistics loss close to 1. The solid line shows the mean loss over all initial conditions, the fine dotted lines are the individual initial conditions; shaded areas show the standard deviation.
@@
For our final learning setup we target a three dimensional scenario with a turbulent channel flow (TCF) setup. This setup contains a periodic channel with no-slip boundaries at $\pm y$ which is driven by a dynamic forcing $\nu\left.\partial\overline{u}/\partial y\right|_{\text{wall}}$ to prevent a loss of energy. The TCF is a well studied scenario that requires high spatial and temporal resolution as well as long simulation times for the turbulence statistics to converge [61]. In this scenario, coarse spatio-temporal resolutions quickly yield incorrect statistics, and hence it is crucial to introduce a form of numerical modeling of the unresolved scales to obtain an accurate solution without excessive runtimes. In line with previous experiments, we train an SGS model in form of a corrector $G_{\theta}$ that is tasked to estimate a correcting force $S_{\theta}$ at a low spatial resolution of $64\times 32\times 32$. For the TCF, $G_{\theta}$ receives the instantaneous velocity and the normalized wall distance $1-|y/\delta|$ as inputs, i.e., $S^{n}_{\theta}=G_{\theta}(\mathrm{\mathbf{u}}^{n},1-|y/\delta|)$. The term $1-|y/\delta|$ is added to inform the network of the grid refinement in regions near the wall. As neural network architecture for $G_{\theta}$ we employ a simple CNN with layers using 8, 64, 64, 32, 16, 8, 4, and 3 filters, each having a kernel size of $3^{3}$. Only the last layer uses a kernel size of $1^{3}$. This gives 198931 trainable parameters in total. ReLU activations are used for all but the last layer. Different from the previous scenarios, the dynamics in the 3D TCF are highly turbulent and matching individual realizations of the flow from simulations at different resolutions no longer provides a physically meaningful learning target. Hence, we instead aim for matching the turbulence statistics via the statistic loss from eq. (13), complemented with a regularization term on the generated forcing.
@@
[…]
We additionally constrain the forcing to the [2,2][-2,2] range to stabilize early training. Aside from the loss terms guiding the model towards producing the desired turbulence statistics, it proved essential to prevent un-physical network outputs, which in our case means violating the incompressibility assumption. To ensure divergence free flow motions, we include the gradient modification from eq. (11) for $S_{\theta}$.
@@
Based on our results from section 4.3, we consider two different approaches for the training of the CNN model: in the first one, we exclude the gradients of the linear solves from the optimization (i.e. only using $J^{\text{none}}$ from eq. (8)). In the second variant, we start in the same way, but include the terms at a later stage of training for fine tuning. The network is trained in four different phases where the warm-up steps are sampled uniformly random from [0,0],[0,12],[0,24],[0,0],~[0,12],~[0,24], and $[0,96]$, respectively. The first phase consists of 6k optimization steps, while the other three include 20k steps. The warm-up is always followed by 12 steps of unrolling, for which we backpropagate through the PICT solver. We denote the result of this initial training as CNN SGS base. We then fine-tune the network with gradients from the advection linear solve, called CNN SGS Adv, for another 20k optimization steps, again with $[0,96]$ steps of warm-up. To ensure a fair comparison, we also continue training with the initial model as CNN SGS for another 20k optimization steps, but without the additional gradient terms. As can be seen from figure 13, the additional gradients from the advection solve only marginally improve the results while roughly doubling the training time. This is consistent with our results from section 4.3 where a rollout of 10 steps did not benefit from the additional gradient terms. The learning task for the TCF models potentially also benefits from the fact that the network still receives feedback for every simulation step via the turbulence statistics loss, making a backpropagated transport of the gradients through the advection unnecessary.
@@
For the no SGS and SMAG variants, the mean streamwise velocity $U^{+}$ deviates noticeably from the reference, although SMAG is closer to the target $\text{Re}_{\tau}$ despite the model having almost no influence in the near wall region. When comparing second order statistics, we can observe how the non-modelled simulation and the simple turbulence model are insufficient to correctly simulate the channel flow at the chosen resolution. Our learned corrector, on the other hand, matches the target statistics very well, and the resulting $\text{Re}_{\tau}=548$ is very close to the target value of 550. Since matching the statistics used in a training loss is expected for a well trained network, we also compare the resulting turbulent energy budgets to corresponding reference values in figure 12. Despite not being trained on these quantities, our model accurately matches the reference budget terms, although slight deviations are noticeable in this evaluation, in particular close to the wall and in the dissipation term. However, it is worth noting that both baselines fail to match any of the budget terms. When investigating the higher order moments, skewness and flatness, the results start to deviate from the reference, with both the CNN SGS and no SGS matching the reference only qualitatively. The CNN SGS also shows deviations near the wall, as shown in B.6.2.
[…]
@@ 5.4 Runtime Performance
[…]
@@ 6 Conclusion
Having established the simulator’s applicability to learning tasks, we showed its efficacy in a number of learning applications. In two challenging 2D settings, we trained stable corrector networks to yield accurate solutions with low-resolution simulations over long rollouts. The learning setup was adapted to three-dimensional turbulent channel flows, enabling the corrector to recover the statistical properties of turbulence over rollout horizons exceeding the training window by orders of magnitude, without requiring direct supervision. Despite the uncertainties inherent to deep learning, such as generalization and the lack of guarantees for unconstrained aspects of the solution, our learned corrector remains remarkably stable, even for different initial conditions. And while it is limited to the geometry seen during training and not necessarily reproduces the details of higher order moments, it still matches the expected energy budgets. It additionally outperforms traditional solvers like OpenFOAM both in terms of accuracy and computational performance. Together, these tests show the efficacy of using learned corrector networks in conjunction with coarse simulations to retain high fidelity behavior in the flows. The availability of the PICT solver as open source [^2] https://github.com/tum-pbs/PICT provides a powerful foundation for advancing learned turbulence modeling for the research community.
All trained models are CNNs, which limits them to regular grid inputs and requires custom padding for multi-block setups. An extension to unstructured meshes is possible via GNNs (graph neural networks), which are typically more expensive. Investing into a differentiable re-gridding to transfer from unstructured to regular meshes is another possibility to run CNNs on unstructured simulations. Our models use the complete velocity fields as input, making the availability of the dense fields a necessity, but allow the model to use more than local information, e.g., the model employed as TCF SGS has a receptive field of $15^{3}$, running on a grid of size $64\times 32\times 32$. If sparse sensors or actors are targeted, local models like those employed in SciMARL [44] could also be trained with our solver. While our solver is implemented with GPU support for increased performance, it is currently limited to a single GPU. As future work, our simulator could be extended to work in multi-GPU setups, and to include multi-grid solvers for the advection and pressure systems, from which we expect further performance benefits. Overall, the presented results are indicative of our solver’s ability to address optimization tasks like initial and boundary value reconstruction, learning for control problems, and, with an extension for differentiable transformations, potentially also for tasks like shape optimization [85]. Moreover, the solver’s full differentiability enables promising opportunities in uncertainty quantification for fluid dynamics through probabilistic and diffusion-based learning frameworks [86, 87].
@@ A.2 PISO Algorithm
which is solved for the pressure $p^{*}$. This pressure is then used to compute the corrected, divergence-free velocity $\mathrm{\mathbf{u}}^{**}$ with
@@
The pressure correction, equations 19, 20 and 21, are repeated twice [5], with an additional indicating the second round of updates. The velocity of the next time-step is then $\mathrm{\mathbf{u}}^{n+1}:=\mathrm{\mathbf{u}}^{***}$.
@@ A.3 Discretization
Set of valid neighbour cells of $P$ $F\in\textbf{F}$
Set of faces between $P$ and F $f\in\textbf{f}$
Set of (virtual) boundary neighbour cells of $P$ $B\in\textbf{B}$
Set of boundary-faces between $P$ and B $b\in\textbf{b}$
@@
Using this notation to reference values in one of the matrices is less straight-forward, since the matrix entries relate to two cells. Thus, with current cell $P$ with index $i$ and neighbor $F\in\textbf{F}$ with index $j$, the entries of a matrix $C$ are referenced as:
@@
Neighbour entry of neighbour cell (its own center) ${\left[C_{F}\right]_{F}}=C_{jj}$